OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x + 4y = -36 4x + 4y = 12

OpenStudy (anonymous):

2 equations 2 unknowns break is between -36 and 4x

OpenStudy (anonymous):

yes 4x+4=-36 4x+4=12

OpenStudy (anonymous):

from the first ewuation 4y= -4x -36

OpenStudy (anonymous):

y= -x -9

OpenStudy (anonymous):

second ewuation : 4y = -4x +12

OpenStudy (anonymous):

y = -x +3

OpenStudy (anonymous):

the two lines have sam slope but different y ontercepts....therefore the lines are parallel and hence no solution